OpenStudy (anonymous):

Suppose f(x)=(2x+1)^10*(3x-1)^7. Find a formula for f'(x).

OpenStudy (gorv):

f'(x) means differentiation of f(x)

OpenStudy (gorv):

find \[\frac{ df(x) }{ dx }\]

OpenStudy (aum):

\[\large f(x) = (2x+1)^{10*(3x-1)^7} \\ \large f(x) = e^{\ln(f(x))} = e^{\ln\{(2x+1)^{10(3x-1)^7}\}} \\ \large f(x) = e^{10(3x-1)^7\ln(2x+1)} \\ \large f'(x) = e^{10(3x-1)^7\ln(2x+1)} * \frac{d}{dx}\{10(3x-1)^7\ln(2x+1)\}~~ =~~ ? \]